# Improving Vision-Language-Action Models via Chain-of-Affordance

### v1

@@ Abstract
Robot foundation models, particularly Vision-Language-Action (VLA) models, have garnered significant attention for their ability to enhance robot policy learning, greatly improving robot’s generalization and robustness. OpenAI’s recent model, O1, showcased impressive capabilities in solving complex problems by utilizing extensive reasoning chains. This prompts an important question: can robot models achieve better performance in multi-task, complex environments by reviewing prior observations and then providing task-specific reasoning to guide action prediction?
In this paper, we introduce Chain-of-Affordance (CoA), a novel approach to scaling robot models by incorporating reasoning in the format of sequential robot affordances to facilitate task completion. Specifically, we prompt the model to consider the following four types of affordances before taking action: (1) object affordance — what object to manipulate and where it is; (2) grasp affordance — the specific object part to grasp; (3) spatial affordance — the optimal space to place the object; and (4) movement affordance — the collision-free path for movement. By integrating this knowledge into the policy model, the robot gains essential context, allowing it to act with increased precision and robustness during inference. Our experiments demonstrate that CoA achieves superior performance than state-of-the-art robot foundation models, such as OpenVLA and Octo. Additionally, CoA shows strong generalization to unseen object poses, identifies free space, and avoids obstacles in novel environments.
@@ 1 Introduction
Recent advancements in Vision-Language-Action (VLA) models have shown that training with internet-scale data can empower end-to-end policy learning models to outperform non-VLA models. However, current approaches often rely heavily on high-level planning or task decomposition by off-the-shelf LLMs or VLMs, limiting models from developing implicit reasoning on their own. OpenAI’s recent O1 model has demonstrated that LLMs can improve performance on complex tasks through extensive reasoning chains. It is expected that this reasoning capability if adapted to robotic models, could enhance their action robustness and generalizability. Yet, the exploration of self-driven reasoning in robotics remains under-explored, highlighting an important frontier for future research.
@@
In this work, we propose Chain-of-Affordance, namely CoA, a novel perspective on generalizing model reasoning at test-time, and leverage such generated reasoning to facilitate the policy learning process. Our model builds upon the DiffusionVLA [43], a state-of-the-art VLM model combining autoregression and diffusion objectives. Our method leverages visual affordance in robot learning, conceptualizing various actions and interactions with objects or the environment that a robot can perform based on visual contexts. We consider four types of affordance that are critical for robots to understand their observational surroundings to interact effectively with objects and achieve tasks in dynamic:
@@
  • Object affordance. When a user provides vague instructions, the robot should be capable of identifying the target object for interaction and its location within the view.
@@
  • Grasp affordance. It involves the robot assessing the object’s most appropriate points or surfaces to enable secure and stable handling.
@@
  • Spatial affordance. The robot needs to identify a set of coordinates that satisfy relationships described in language, such as free space for placement.
@@
  • Movement affordance. Identifying a trajectory for the robot to move without collision is crucial in the real world to avoid catastrophic damage.
@@
These four affordances form a sequential chain, requiring the robot to possess prior knowledge at each step to advance to the next. Specifically, at test time, the robot must first understand ”what to manipulate and where the object is located.” Next, it determines ”how to grasp the object” and finally, ”where to place it,” all while following a trajectory that ensures safe task completion. During inference, the affordance chain is progressively generated as the action state evolves, avoiding unnecessary computational costs associated with outputting extensive language. We introduce two formats for chain-of-affordance reasoning: text-based and image-based chain-of-affordance prompting. Our model’s architecture natively supports text prompting integration, but image prompting requires adaptation. To address this, we developed an image affordance injection module that seamlessly integrates visual affordances into the policy network. Once generated by the model, this affordance knowledge is reused in policy learning, enabling the model to produce robust and generalizable actions.
@@
We conduct extensive experiments on both simulated benchmarks and real-world tasks. Specifically, on the LIBERO [27] benchmark, our proposed CoA outperforms several other state-of-the-art approaches, including the Diffusion Policy [9], Octo [33], and OpenVLA [21]. Due to the substantial sim-to-real gap, many existing studies also evaluate real robots. We set up seven real-world robot tasks, including long-horizon and challenging tasks such as serving tea, cleaning garbage, and wiping tables. We perform multi-task learning on these real-world tasks and demonstrate that the VLA model, enhanced by chain-of-affordance reasoning, successfully handles most of these complex tasks with a high average success rate. Furthermore, CoA exhibits strong generalization capabilities to handle unseen object poses, adjust motion when obstacles appear, and identify free space for placement.
@@ 2 Related Works
Reasoning for language and control. Prompting large language models (LLMs) with ”think step-by-step” [42] has significantly advanced their ability to solve complex tasks. Numerous approaches [3, 50] have since been developed to encourage deeper reasoning in LLMs, such as Tree-of-Thought [45] and Chain-of-Code [23], establishing this as a standard practice in language modeling. Recent research has leveraged LLMs and vision-language models (VLMs) [15, 13, 25, 10, 16, 14] as high-level planners in robotics, often using fine-tuned, open-source models or closed-source LLMs alongside policy networks for low-level task execution. These studies illustrate that detailed reasoning can enhance low-level control. ECoT [48] introduces a reasoning strategy for VLMs that includes task decomposition, subtask descriptions, fine-grained movement instructions, gripper positioning, and object tracking on the table. However, the extensive output increases computational costs. Our work proposes a more concise and efficient reasoning approach, improving low-level control while reducing computational demands.
@@ 3 Preliminary on Vision-Language-Action Models
Our work builds upon Vision-Language-Action (VLAs) as the backbone for our chain-of-affordance policies. VLAs employ a straightforward policy-learning approach: beginning with a pre-trained vision-language model, they fine-tune it to predict the next robot action $a$ based on the current image observation $I$, task instruction $T$, and reasoning $r$. There are two types of VLAs: autoregressive VLAs [7, 34, 1, 48, 21] and diffusion-based VLAs [44, 24, 6, 43]. The former uses discrete action tokens within the vision-language model’s vocabulary, enabling action generation similar to language modeling through next-token prediction. The latter leverages a policy head [5, 4, 26], such as a diffusion policy [44] or flow matching [6], to output continuous robot actions.
In this work, we employ the recently released DiffusionVLA [43] model, which integrates the Qwen2-VL [41] vision-language model with a diffusion model head for action prediction. Specifically, the model outputs language through the VLM backbone and the language embeddings are subsequently injected into the policy model to enhance policy learning. This operation allows the model to reuse reasoning information while achieving faster inference than ECoT [48]. In the following sections, we will discuss how we improve this VLA by enabling it to reason through robot affordances before selecting an action.
@@ 4 Methodology
Figure 1: An example of the chain of affordances in the Pour Tea task is shown here. The first row presents the text-based chain-of-affordance prompts, while the second row shows the image-based chain-of-affordance prompts.
@@ 4.1 Definition of Chain-of-Affordance
Object affordance: Object affordance enables the robot to determine what object to interact with and where it is located, particularly when a user query lacks explicit instructions. This foundational knowledge allows the robot to discern the object it should manipulate and localize it within its environment. In our setup, we require the robot to identify the object by name in natural language and predict its location in an image using a bounding box format. This identification provides the base knowledge for the robot’s decision-making and interaction processes.
@@
Spatial affordance: Spatial affordance is about spatial relations as an indicator of a model’s ability to understand a 3D world. This affordance supports tasks like identifying free spaces for object placement or navigation. For instance, RoboPoint [47] locates free space, and SpatialVLM [8] predicts spatial relations quantitatively and qualitatively. In our work, we define an actionable destination for the robot, such as vacant space on a plate for placing objects or a clear path for cleaning. We represent spatial affordance as a set of 2D coordinates to indicate feasible areas for action.
@@
Updating affordance chain. During both the training and inference stages, only a subset of affordance is involved based on the robot and scene state. Specifically, the model first predicts object affordance and grasp affordance, which tells the model what and how to manipulate. After the object is captured by the robot, it predicts the spatial affordance. The movement affordance is predicted throughout the whole period of action. By updating the affordance chain in real-time, we ensure minimum computational overhead is introduced.
Notably, the language generation is optimized through next token prediction, and the action generation is updated via diffusion model objective. This technique is similar to visual grounding works. We utilized the pre-trained checkpoint of DiffusionVLA to initialize our model. All constructed data are co-trained.
4.2 Formatting of Chain-of-Affordance
In this section, we introduce two formats for representing the chain of affordance: one in text format and another in image format. The text-based format provides a structured, linguistic approach to affordance representation, while the image-based format offers a visual perspective that can enhance the interpretability of affordance cues in visually complex environments. Additionally, we describe a new module specifically designed to integrate the image-based chain of affordance prompts into the policy model, thereby enabling the model to leverage both visual and linguistic cues for action generation.
Text-based chain-of-affordance prompting. Natural language is the most common way to represent visual affordances. For example, in object affordance, a set of coordinates can specify the location of objects, while spatial affordance can be conveyed through another set of coordinates. Figure 1 (top) provides an example of using natural language to denote various affordances. In practice, we avoid using a fixed language template to describe affordances. Instead, we use ChatGPT to enrich language expression by varying formats, similar to data augmentation in natural language processing. This approach helps the model retain its conversational capabilities. Notably, our base model is inherently capable of reasoning in natural language, so we adhere to its architecture to process affordances in text format to the policy model.
Image-based chain-of-affordance prompting. Inspired by methods that enhance model performance through visual prompts, we introduce visual affordance augmentation, which supplements natural language descriptions with direct visual cues. This approach enables the model to interpret and utilize affordance information more effectively.
Specifically, we overlay coordinates or point trajectories directly onto the robot’s previous observation frame, serving as a form of chain-of-affordance prompting that visually guides the model. This overlay provides a clear, interpretable structure, linking objects and their potential actions within the scene, and helps bridge the gap between descriptive language and practical visual information. Examples of this approach are shown in Figure 1 (bottom), where affordances are visually layered onto the image.
For movement affordances, which are applicable across all scenarios, we use a thinner point trajectory. This design choice prevents these trajectories from drawing excessive attention. Additionally, we employ thicker and lighter colors to denote key elements such as the object’s bounding box, grasping points, and spatial affordances. These visual distinctions ensure that each affordance type is easy to identify and interpret at a glance, enabling the model to parse visual information effectively.
Injecting image-based chain-of-affordance prompting. Unlike text-based chain-of-affordance prompting, the base model does not natively support images as intermediate inputs to the policy network. To overcome this limitation, we propose an image chain-of-affordance injection module that incorporates visual knowledge directly into the policy model. Our approach involves augmenting the robot’s current observation with visual affordances, allowing the model to retain and utilize affordance-based cues directly from images. Specifically, we use the pre-trained visual encoder from base VLA to extract detailed feature representations from the image, breaking it down into multiple patches to capture a finer level of detail and spatial context. These visual tokens are concatenated and passed through Multi-Layer Perceptron (MLP) layers to reduce the channel dimensions. To inject the image chain-of-affordance prompting seamlessly, we leverage Feature-wise Linear Modulation (FiLM) [36], which enables adaptive conditioning of the model based on visual context. This branch is similar to the reasoning injection module in the DiffusionVLA [43]. By aligning image affordances with FiLM, we ensure that the model can interpret and respond to visual cues in its decision-making process effectively. The image-based affordance injection allows the model to process richer contextual information.
@@ 4.2 Formatting of Chain-of-Affordance
4.3 Generating Chain-of-Affordance Data
Our pipeline begins with GPT-4o [35], which generates a detailed description of the scene and identifies relevant entities from the language instructions. This allows us to create contextually rich, entity-specific affordances tailored to the task at hand. Using these entities, we leverage Grounding DINOv2 [29] and SAM [22, 37] to produce bounding boxes around each identified object within the scene. SAM initially provides masks, which we convert into bounding boxes, and we finalize these by taking out the intersection over union (IoU) between the outputs of Grounding DINOv2 and SAM. This IoU-based refinement ensures that bounding boxes are accurately aligned with object contours. At this stage, we also capture the gripper’s position for subsequent affordance calculations.
To represent spatial affordance, we integrate RoboPoint [47], a state-of-the-art model that predicts spatial affordances directly within the image. Additionally, we prompt GPT-4o to annotate spatial points based on the scene context. The spatial predictions from RoboPoint and GPT-4o are then combined, after which we cluster these points to form a coherent representation, eliminating any outliers to maintain accuracy. This process captures spatial affordances that are not only visually aligned but contextually relevant.
For capturing movement trajectories, we employ CoTracker [18, 19], an advanced transformer-based tracking model. CoTracker enables us to follow the robot gripper’s path, recording its movement through the scene and gathering essential trajectory data. This movement data provides insight into how the robot interacts dynamically with the environment, adding temporal dimensions to our affordance representation. The combination of spatial affordances, grasping points, and movement trajectories enables us to model a rich, multi-faceted affordance landscape for each scenario. This automated, tool-assisted pipeline produces a detailed and diverse set of affordance annotations, significantly reducing the need for manual labeling.
5 Experiments
The objective of our experimental evaluations is to assess the efficacy of CoA as a reasoning framework that augments the policy-learning capabilities of a baseline robot foundation model. Specifically, we seek to address the following questions:
  • Does the robot foundation model demonstrate improved performance with the integration of CoA in real-world experiments?
  • What is the performance of CoA in simulation benchmarks?
  • How effectively does CoA generalize to challenging scenarios?
5.1 Evaluation on Real Robot
Experimental setup. We conduct evaluations using the 6-DoF Franka robot arm, a widely adopted setup for assessing generalizable robot policies. Our setup includes two third-person cameras (ZED cameras) positioned on either side of the robot arm, as well as an egocentric camera (Realsense D435i) mounted on the wrist. We design seven challenging tasks, such as pouring tea and cleaning the table, incorporating both short-horizon and long-horizon tasks to test the adaptability of our approach. We illustrate our robot setup and tasks with examples in Figure 2. Detailed descriptions of each task and the experimental setup are provided in the Appendix.
Train data. We utilize the Droid dataset [20] as an external data source. Specifically, since not all samples in Droid contain annotations for language instructions, we filter out these samples, resulting in a total of 39K trajectories. We then apply our pipeline to generate synthetic chain-of-affordance data for Droid and use this data for pre-training. Following this, we post-train the model on task-specific data, collecting 692 trajectories for seven tasks in total.

### v2

# CoA-VLA: Improving Vision-Language-Action Models via Visual-Textual Chain-of-Affordance

@@ Abstract
Robot foundation models, particularly Vision-Language-Action (VLA) models, have garnered significant attention for their ability to enhance robot policy learning, greatly improving robot’s generalization and robustness. OpenAI’s recent model, O1, showcased impressive capabilities in solving complex problems by utilizing extensive reasoning chains. This prompts an important question: can robot models achieve better performance in multi-task, complex environments by reviewing prior observations and then providing task-specific reasoning to guide action prediction? In this paper, we introduce Chain-of-Affordance (CoA-VLA), a novel approach to scaling robot models by incorporating reasoning in the format of sequential robot affordances to facilitate task completion. Specifically, we prompt the model to consider the following four types of affordances before taking action: (1) object affordance — what object to manipulate and where it is; (2) grasp affordance — the specific object part to grasp; (3) spatial affordance — the optimal space to place the object; and (4) movement affordance — the collision-free path for movement. We further transform each affordance into two prompting formats: visual affordance and textual affordance. We introduce a novel vision-language co-injection module that integrates this knowledge into the policy network. This allows the robot to leverage essential contextual information during action inference, resulting in improved precision and robustness. Our experiments demonstrate that CoA-VLA outperforms state-of-the-art robot foundation models, including OpenVLA and Octo, on a variety of tasks. Furthermore, CoA-VLA exhibits strong generalization capabilities, including recognizing unseen object poses, identifying free space, and avoiding obstacles in novel environments.
Figure 1: This figure illustrates the overall framework of our CoA-VLA model, which empowers vision-language-action models with chain-of-thought reasoning capabilities for generalizable visuomotor policy learning. We achieve this by designing four distinct types of affordance and introducing a novel visual-text co-injection method to integrate this knowledge into the decision-making process.
@@ 1 Introduction
Recent advancements in Vision-Language-Action (VLA) models have shown that training with internet-scale data can empower end-to-end policy learning models to outperform non-VLA models. However, current approaches often rely heavily on high-level planning or task decomposition by off-the-shelf Large language models(LLMs) or Vision language models (VLMs), limiting models from developing implicit reasoning on their own. OpenAI’s recent O1 model has demonstrated that LLMs can improve performance on complex tasks through extensive reasoning chains. If this reasoning capability can be applied to the control model, it could enhance their action robustness and generalizability. Yet, the exploration of self-driven reasoning in robotics remains under-explored, highlighting an important frontier for future research.
@@
In this work, we propose Chain-of-Affordance, namely CoA-VLA, a novel perspective on generalizing model reasoning at test-time, and leverage such generated reasoning to facilitate the policy learning process. Our model builds upon the DiffusionVLA [46], a state-of-the-art VLM model combining autoregression and diffusion objectives. Our method leverages visual affordance in robot learning, conceptualizing various actions and interactions with objects or the environment that a robot can perform based on visual contexts. We consider four types of affordance that are critical for robots to understand their observational surroundings to interact effectively with objects and achieve tasks in dynamic:
@@
1) Object affordance. When a user provides vague instructions, the robot should be capable of identifying the target object for interaction and its location within the view.
@@
2) Grasp affordance. It involves the robot assessing the object’s most appropriate points or surfaces to enable secure and stable handling.
@@
3) Spatial affordance. The robot needs to identify a set of coordinates that satisfy relationships described in language, such as free space for placement.
@@
4) Movement affordance. Identifying a trajectory for the robot to move without collision is crucial in the real world to avoid catastrophic damage.
@@
These four affordances form a sequential chain, requiring the robot to possess prior knowledge at each step to advance to the next. Specifically, at test time, the robot must first understand “what to manipulate and where the object is located.” Next, it determines “how to grasp the object and finally”, “where to place it,” all while following a trajectory that ensures safe task completion. During inference, the affordance chain is progressively generated as the action state evolves, avoiding unnecessary computational costs associated with outputting extensive language. We introduce two formats for chain-of-affordance reasoning: text-based and image-based chain-of-affordance prompting. Our model’s architecture natively supports text prompting integration, but image prompting requires adaptation. To address this, we developed an image affordance injection module that seamlessly integrates visual affordances into the policy network. Once generated by the model, this affordance knowledge is reused in policy learning, enabling the model to produce robust and generalizable actions.
@@
We conduct extensive experiments on both simulated benchmarks and real-world tasks. Specifically, on the LIBERO [30] benchmark, our proposed CoA-VLA outperforms several other state-of-the-art approaches, including the Diffusion Policy [10], Octo [36], and OpenVLA [22]. Due to the substantial sim-to-real gap, many existing studies also evaluate real robots. We set up seven real-world robot tasks, including long-horizon and challenging tasks such as serving tea, cleaning garbage, and wiping tables. We perform multi-task learning on these real-world tasks and demonstrate that the VLA model, enhanced by chain-of-affordance reasoning, successfully handles most of these complex tasks with a high average success rate. Furthermore, CoA-VLA exhibits strong generalization capabilities to handle unseen object poses, adjust motion when obstacles appear, and identify free space for placement.
The core contribution of our work lies in a novel framework that enhances vision-language-action (VLA) models by integrating affordance-aware reasoning. Central to this framework is a new module designed to strategically infuse affordance knowledge into policy learning, enabling robots to better ground actions in physical and contextual understanding. While our approach draws inspiration from prior studies on affordance, the synthesis of these concepts with VLA systems represents a significant advancement. Specifically, our framework uniquely combines these components to forge more robust, interpretable, and generalizable robotic policies, addressing critical limitations in existing VLA systems that often overlook the role of object affordances in decision-making.
@@ 2 Related Works
Reasoning for language and control. Prompting large language models (LLMs) with ”think step-by-step” [45] has significantly advanced their ability to solve complex tasks. Numerous approaches [3, 55] have since been developed to encourage deeper reasoning in LLMs, such as Tree-of-Thought [49] and Chain-of-Code [24], establishing this as a standard practice in language modeling. Recent research has leveraged LLMs and vision-language models (VLMs) [16, 14, 27, 11, 17, 15, 17] as high-level planners in robotics, often using fine-tuned, open-source models or closed-source LLMs alongside policy networks for low-level task execution. These studies illustrate that detailed reasoning can enhance low-level control. ECoT [52] introduces a reasoning strategy for VLMs that includes task decomposition, subtask descriptions, fine-grained movement instructions, gripper positioning, and object tracking on the table. CoT-VLA [54] generates sub-goa for autoregressive VLA to guide action. Unlike existing approaches, our work introduces an affordance taxonomy (categorized into four types) to unify textual and visual affordance representations in robot learning. Coupled with a dynamic selection mechanism, this taxonomy enables adaptive policy training, where task-relevant affordances are prioritized at each timestep. The resultant framework achieves computationally efficient reasoning while retaining robustness to environmental ambiguities.
@@ 3 Preliminary on Vision-Language-Action Models
Our work builds upon Vision-Language-Action (VLAs) as the backbone for our chain-of-affordance policies. VLAs employ a straightforward policy-learning approach: beginning with a pre-trained vision-language model, they fine-tune it to predict the next robot action $a$ based on the current image observation $I$, task instruction $T$, and reasoning $r$. There are two types of VLAs: autoregressive VLAs [8, 37, 1, 52, 22, 39, 29] and diffusionVLAs [47, 26, 7, 46]. The former uses discrete action tokens within the vision-language model’s vocabulary, enabling action generation similar to language modeling through next-token prediction. The latter leverages a policy head [6, 5, 28], such as a diffusion policy [47, 25] or flow matching [7], to output continuous robot actions.
In this work, we employ the recently released DiVLA [46] model, which integrates the Qwen2-VL [44] vision-language model with a diffusion model head for action prediction. In the following sections, we will discuss how we improve this VLA by enabling it to reason through robot affordances before selecting an action.
@@ 4 Methodology
Figure 2: An example of the chain-of-affordance for the PourTea task. The first row presents the text affordance and the second row shows the visual affordance. By employing a dynamic affordance selection mechanism, our method avoids generating redundant affordances at every timestep.
@@ 4.1 Definition of Chain-of-Affordance
Object affordance: Object affordance equips robots with the foundational ability to autonomously determine which object to interact with and where it is located, particularly in scenarios where user instructions lack explicit spatial or semantic details. This capability bridges the gap between ambiguous queries (e.g., “Pour the drink”) and actionable execution by enabling the robot to: 1) Identify the target object through natural language grounding (e.g., resolving “drink” to “teapot”), 2) Localize the object within its environment using pixel-aligned bounding box predictions. In our framework, object affordance is operationalized through two tightly coupled components: 1) Semantic identification: Resolving object names from free-form language input, and 2) Spatial grounding: Predicting the object’s 2D location via visual scene understanding. By integrating these capabilities, the robot establishes a contextual foundation for downstream decision-making, ensuring interactions are both intention-aware (aligned with user goals) and environmentally grounded (physically feasible).
@@
Spatial affordance: Spatial affordance centers on a robot’s ability to interpret and reason about spatial relationships within 3D environments, enabling tasks such as identifying collision-free regions for object placement or navigation. For instance, RoboPoint [51] locates free space, and SpatialVLM [9] predicts spatial relations quantitatively and qualitatively. In our framework, spatial affordance is operationalized as actionable destinations—discrete 2D coordinates representing feasible interaction zones.
@@
Dynamic affordance selection. Our proposed method formulates affordances as combined text and visual prompts. While this approach offers benefits, it introduces additional computation at test time, potentially slowing down the algorithm. We observe that it’s unnecessary to utilize all available affordances during testing. For example, once an object is picked up, predicting its object affordance and the grasp affordance becomes redundant. The specific affordances required in a given sub-step depend on the task’s progress and the environment. Therefore, we implement dynamic affordance selection, adaptively choosing the necessary affordances at both training and test times to reduce computational cost. Several methods can achieve this, such as gradient-based selection [48]. Our approach prioritizes simplicity by leveraging proprioception. Proprioception refers to information about the robot’s state, including joint angles and other movement data. We transform this proprioceptive data into a single token and concatenate it with the visual token before feeding it into the large language model. Training on large-scale annotated datasets like Droid [21] enables our model to intelligently select relevant affordances at each time step. This is straightforward for the model to learn. For instance, if the proprioceptive state indicates a partially closed gripper and the wrist-mounted camera detects an object, the model can infer that the object and grasp affordances are likely unnecessary. Instead, it can focus on the movement affordance to guide the action trajectory and the spatial affordance to determine a suitable placement location. We found this strategy to be simple and useful, reducing the model’s computational cost without hurting performance.
@@
4.2 Formatting of Visual-Textual Chain-of-Affordance
In this section, we present a multimodal chain-of-affordance framework that formalizes affordance reasoning through two complementary representations: (1) textual affordance, which provides a structured, language-driven representation of actionable possibilities (e.g., “graspable handle” or “navigable path”), and (2) visual affordance, which delivers a spatially grounded, scene-aware perspective via pixel-aligned cues. While the textual format enables explicit semantic reasoning, the visual modality enhances interpretability in visually dense or ambiguous environments by grounding affordances in observable scene geometry.
To unify these modalities, we introduce a visual-language co-injection module that dynamically aligns and integrates visual affordance prompts in a shared parameter space. This module bridges the gap between abstract language-based reasoning and pixel-level visual context, enabling the policy model to synergistically leverage both modalities for robust, context-aware action generation.
Textual affordance. Natural language serves as the predominant modality for representing visual affordances due to its semantic richness and compatibility with human-AI interaction. For instance, object affordances can be encoded as coordinate pairs (e.g., bounding boxes localizing a ”graspable cup”), while spatial affordances are expressed as coordinate pairs defining navigable regions (e.g., ”reachable shelf area”). Figure 2 (top) illustrates how diverse affordances (e.g., ”movable,” ”stackable”) are mapped to natural language descriptors.To avoid rigid linguistic templates that might bias policy learning, we employ ChatGPT to dynamically paraphrase affordance descriptions with varied syntactic structures and vocabulary (e.g., alternating between ”placeable surface” and ”flat area for stacking”). This strategy, analogous to data augmentation in NLP, enhances the model’s robustness to linguistic variability while preserving its conversational fluency.
Visual affordance. Inspired by visual prompting methods that enhance model interpretability, we propose visual affordance augmentation—a technique that augments natural language affordance descriptions with direct, pixel-aligned visual cues. This approach encodes affordances by overlaying coordinate markers (e.g., bounding boxes, interaction points) or motion trajectories onto the robot’s historical observation frames (Figure 2, bottom). These annotations act as chain-of-affordance prompts, visually grounding the model’s understanding of actionable properties like ”graspable” or ”navigable” within the scene’s geometry. By embedding affordances directly into the visual input, we create an explicit structure that bridges the gap between abstract language and actionable visual context.For movement affordances, we employ thin, low-saliency trajectories to avoid overshadowing critical scene elements while guiding motion planning. Conversely, key interaction zones — such as grasping points, object bounding boxes, and spatial affordances — are rendered with thicker, high-contrast overlays (e.g., semi-transparent colors) to ensure visual salience without occluding environmental context. This hierarchical visual encoding distinguishes affordance types at a glance, enabling the model to prioritize task-relevant cues while maintaining computational efficiency.
Visual-textual co-injection. To integrate visual and textual affordances into action generation, we design a unified embedding pipeline that seamlessly fuses both modalities into the diffusion policy. For textual affordances, we use the last embedding from the VLM models and add an MLP layer to tokenize it. Concurrently, visual affordances are processed into patch tokens via a pretrained Vision Transformer (ViT-Small). The fused tokens are then processed by two standard Transformer blocks, balancing computational efficiency with sufficient expressive power for cross-modal reasoning. The encoder’s output embeddings are projected into the diffusion model using FiLM conditioning layers, inspired by MT-ACT [4]. This conditioning dynamically modulates the diffusion process by injecting affordance-aware features, enabling the policy to generate actions that respect both spatial constraints and semantic intent. By unifying affordance modalities early in the pipeline, our method retains the computational efficiency of standard diffusion frameworks while enhancing robustness. The FiLM-based conditioning acts as a bottleneck, distilling only the most salient affordance cues to guide action generation, thereby avoiding overfitting to redundant visual or linguistic signals.